The following information is available for Rodriguez Industries:

Direct labor $ 84,000
Total current manufacturing costs 385,000

Manufacturing overhead is applied to production at 150 percent of direct labor cost. Determine the amount of direct materials used in production.

Job 31 has a direct materials cost of $280 and a total manufacturing cost of $820. Overhead is applied to jobs at a rate of 200 percent of direct labor cost.

Use the relationships among total manufacturing costs, conversion cost, and prime cost to determine the following costs for Job 31:

1conversion cost

2direct labor cost

3 manufactring overhead

4prime cost

Fairfield Company’s payroll costs for the most recent month are summarized here:

Item Description Total Cost
Hourly labor wages 810 hours @ $16 per hour
200 hours for Job 101 = $ 3,200
330 hours for Job 102 = 5,280
280 hours for Job 103 = 4,480 $ 12,960
Factory supervision 3,950
Production engineer 6,300
Factory janitorial work 1,800
Selling, general, and administrative salaries 9,700
Total payroll costs $ 34,710

1. Calculate how much of the labor costs would be added to the following accounts:

a. Work in Process Inventory
b. Manufacturing Overhead
c. Selling, General, and Administrative Expenses
Answer #1

Solution 1:

Manufacturing overhead applied = Direct labor*150% = 84000*150% = $126,000

direct materials used in production = Total current manufacturing costs - Direct labor cost - Manufacturing overhead applied

= $385000 - $84000 - $126000 = $175,000

Solution 2:

Conversion cost = total manufacturing cost  - Direct materials cost = $820 - $280 = $540

Direct Labor Cost:

Direct Labor + Manufacturing overhead = Conversion cost

Direct labor + 200% of Direct labor cost = $540

300% Direct labor Cost = $540

Direct labor Cost = $540 / 300% = $180

Manufacturing overhead = $180*200% = $360

Prime Cost = Direct materials + Direct labor = $280 + $180 = $460

Solution 3:

Work in Process Inventory = Direct labor wages = $12,960

Manufacturing Overhead = Factory supervision + Production engineer + Factory janitorial work

= 3950 +6300+ 1800 = $12,050

Selling, General, and Administrative Expenses = $9,700